OpenStudy (anonymous):

a car is moving on a straight road with a speed of 20m/s. At t=0 the driver applies brakes after watching an obstacle 150m ahead. after application of brakes the car retards with 2m/s^2. Find the position of the car from the obstacle at t=15

OpenStudy (anonymous):

i got the answer do u want to answer?

OpenStudy (anonymous):

still reading the question..

OpenStudy (stormfire1):

\[d=V_it+\frac{1}{2}at^2\]\[d=20 m/s(15)+\frac{1}{2}(-2m/s^2)(15)^2\]\[d=75m\]So the car is still 75m from the obstacle after 15 seconds.

OpenStudy (stormfire1):

Is that what you got?

OpenStudy (anonymous):

No. I was asking cause i was getting the wrong answer and doing the same mistake. instead of 2nd equation use 3rd equation of motion then u will ge the answer.

OpenStudy (stormfire1):

yea...I was looking at it and it doesn't seem right at all...let me look again

OpenStudy (anonymous):

the car will stop after 10 seconds of breaking right ? ?

OpenStudy (stormfire1):

yea

OpenStudy (stormfire1):

so yea...I shoulda used 10s

OpenStudy (anonymous):

no 15

OpenStudy (stormfire1):

Well, the car isn't going to start going backwards from braking :P

OpenStudy (anonymous):

after the 1st second the speed will be 18 after second 16 and so on

OpenStudy (anonymous):

use 3rd equation of motion \[v^2-u^2=2as\]

OpenStudy (anonymous):

Initial Velocity (u)=20m/s Acceleration (a) =-2m/s^2 Final velocity v =0 Time(t)=15s #rd equation of motion \[v^2-u^2=2as\] \[0^2-(-20)^2]=2*(-2)s\] \[-400=-4s\] \[-400/-4=s\] \[100=s\]

OpenStudy (anonymous):

3rd equation*

OpenStudy (stormfire1):

Yea, you get 100s when you use the first equation with 10s also :)

OpenStudy (stormfire1):

I just didn't consider that I had to drop off the last 5 seconds since the car was stopped by then.

OpenStudy (anonymous):

yeah but we can't change the given values

OpenStudy (anonymous):

u r also right but the values are given

OpenStudy (stormfire1):

You're not changing anything...it's part of solving the problem in my opinion. If the question had stated 2000seconds...the reality is that the car is still stopped after 10s.

OpenStudy (anonymous):

exactly

OpenStudy (stormfire1):

Besides, when you did your equation, you *assumed* Vf was 0 :)

OpenStudy (anonymous):

u are also right man. consider this <======

OpenStudy (anonymous):

its not assumption if the car stops it means v= 0

OpenStudy (stormfire1):

Did the question state that the car stopped?

OpenStudy (stormfire1):

No :)

OpenStudy (anonymous):

yes

OpenStudy (stormfire1):

What part? It says it's slowing down at 2 m/s...

OpenStudy (anonymous):

no sorry :p

OpenStudy (stormfire1):

So YOU CHANGED THE PROBLEM!! just kidding

OpenStudy (anonymous):

lol

OpenStudy (stormfire1):

Here...let's graph it too so we can beat this problem to death :)